OpenStudy (fanduekisses):

So I did this math problem but I feel like I might've made a mistake because my answer is just weird

OpenStudy (fanduekisses):

\[14\int\limits_{0}^{1}x^3\sqrt(1-x^2)dx\]

OpenStudy (fanduekisses):

I used trig substitution:

OpenStudy (fanduekisses):

\[\sin \theta =\frac{ x }{ 1 }\] \[dx=\cos \theta dt \theta\]

OpenStudy (fanduekisses):

\[14\int\limits_{0}^{1}\sin^3 \theta \sqrt(1-\sin^2)* \cos \theta d \theta\]

OpenStudy (fanduekisses):

Simplified: \[14\int\limits_{0}^{1}\sin^3 \theta \cos^2 \theta d \theta\]

OpenStudy (fanduekisses):

split the sin^3 theta to be able to do U substitution

OpenStudy (fanduekisses):

\[14\int\limits_{0}^{1}\sin^2 \theta \cos^2 \theta \sin \theta d \theta\]

OpenStudy (fanduekisses):

\[u= \cos \theta\] \[du= - \sin \theta d \theta\]

OpenStudy (fanduekisses):

How's it looking so far?

OpenStudy (.sammi.):

Gimme a min

OpenStudy (.sammi.):

How about \[u=x^2~~~du=2xdx\]

OpenStudy (.sammi.):

Then, \[14 \int\limits^1_0 x^3 \sqrt{1-x^2} \, dx \\ \\ 2 \int\limits^1_0 u \sqrt{1-u} \, du\]

OpenStudy (.sammi.):

Get it?

OpenStudy (.sammi.):

I'm just splitting the \(x^3\) into \(x^2\) and \(x\), and from the derivation letting \(u=x^2\), \(du=2xdx\), Then \(\frac{du}{2}=xdx\) \[14 \int\limits\limits^1_0 x^3 \sqrt{1-x^2} \, dx \\ \\ 14 \int\limits\limits^1_0 x^2 \sqrt{1-x^2} (x)\, dx \\ \\ 2 \int\limits\limits^1_0 u \sqrt{1-u} \, du \]

zepdrix (zepdrix):

Your u-sub looks great :) But you have too many sines floating around. You'll need to apply your Pythagorean Identity, \(\large\rm 14\int \color{orangered}{sin^2\theta}cos^2\theta~sin\theta~d \theta\) \(\large\rm 14\int \color{orangered}{(1-cos^2\theta)}cos^2\theta~sin\theta~d \theta\) And at this point, we should remove the limits of integration. They are not limits of theta, they were applying to x.

OpenStudy (.sammi.):

The 2 is actually 7 btw, I made a typo

OpenStudy (fanduekisses):

yeah that is what I did let me show you

OpenStudy (fanduekisses):

\[14\int\limits_{0}^{1}(1-\cos^2 \theta)\cos^2 \theta \sin^2 \theta d \theta\]

OpenStudy (.sammi.):

You don't have to use trig-sub, from here \[7 \int\limits\limits\limits^1_0 u \sqrt{1-u} \, du\] Let \(h=1-u\), \(dh=-du\) \[7 \int\limits\limits\limits^1_0 (h-1)\sqrt{h}\, dh\] Integrate it now

OpenStudy (fanduekisses):

\[u= \cos \theta\] \[du= -\sin \theta \]

zepdrix (zepdrix):

But that's the method he/she choose to use.

zepdrix (zepdrix):

chose*

OpenStudy (fanduekisses):

\[-14\int\limits_{1}^{\cos(1)}(1-u^2)u^2 du\]

zepdrix (zepdrix):

Ooo your limits look strange.. let's just leave them off for now.

OpenStudy (fanduekisses):

\[-14\int\limits_{1}^{\cos(1)}u^2-u^4du= -14[\frac{ u^3 }{ 3 }-\frac{ u^5 }{ 5 }]\]

OpenStudy (fanduekisses):

ohh yeah that part confused me a bit

zepdrix (zepdrix):

Let's factor a 15 out, should clean up the fractions, and distribute the negative,\[\large\rm =\frac{14}{15}(3u^5-5u^3)\]Then we'll undo our u-substitution as a first step of unwinding things, ya?

zepdrix (zepdrix):

Factored a 1/15 out* not 15, blah

OpenStudy (fanduekisses):

oh ok so we have \[-\frac{ 14 }{ 15 }[3\cos^5 \theta - 5\cos^3 \theta]\]

zepdrix (zepdrix):

No more negative in front, if we have the 5th power term leading.

OpenStudy (fanduekisses):

ohh right I see now

zepdrix (zepdrix):

Ummm to get back to x is another thing... we can do fancy triangle business... but, since we were given limits for x, we probably can avoid doing that.

zepdrix (zepdrix):

\(\large\rm x=sin\theta\) x=0: \(\rm 0=sin\theta,\qquad\qquad\theta=?\) x=1: \(\rm 1=sin\theta,\qquad\qquad\theta=?\)

OpenStudy (fanduekisses):

x= 0, theta= arc sin(0) x=1 theta= arc sin(1)

OpenStudy (fanduekisses):

x= 0, theta= 0 x=1 theta= pi/2

zepdrix (zepdrix):

\[\large\rm \frac{ 14 }{ 15 }[3\cos^5 \theta - 5\cos^3 \theta]_0^{\pi/2}\]Ok good.

OpenStudy (fanduekisses):

I get 28/15

OpenStudy (fanduekisses):

Thanks! :D